The graph of a polynomial has x-intercepts at 1, -2, and 10 and passes through the point (2,-32). Determine the equation of the function in factored form.
Found 2 solutions by roseo, solver91311:
Answer by roseo(33)

Because 1,-2, and 10 are x-intercepts or zeros of the polynomial then (x-1), (x+2), and (x-10) are factors of the polynomial.
So we know that
y = a(x-1)(x+2)(x-10) is our polynomial and all we have to do is find a and the point (2,-32) gives us an x and y to put in the function to determine a.
-32 = a (2-1)(2+2)(2-10)
-32 = a (1)(4)(-8)
-32 = -32a
therefore a = 1
So the polynomial y = 1(x-1)(x+2)(x-10) will pass through (2,-32)

Answer by solver91311(24713)



The factored form of all polynomial functions that have -intercepts at 1, -2, and 10 is:



However, we want the particular function such that when , so substitute:



and solve for







Hence:




John